OpenStudy (firejay5):

If f(x) = x^2, g(x) = 5x, and h(x) = x + 4, find each value. Show work and explain please for both problems 14. [fo(hog)](-1) 15. [fo(goh)](4)

OpenStudy (firejay5):

@Mertsj

jimthompson5910 (jim_thompson5910):

[fo(hog)](-1) is the same as saying f [ h(g(-1)) ]

jimthompson5910 (jim_thompson5910):

so step 1) find g(-1) step 2) plug the result from step 1 into h(x) step 3) plug the result of step 2 into f(x) this will give you your final answer for #14

jimthompson5910 (jim_thompson5910):

# 15 is the exact same as #14 but now you're using the functions in a different order and you're using a different initial input

OpenStudy (firejay5):

Can you do #14 with me and let me do #15 on my own? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

ok first step is to find g(-1)

jimthompson5910 (jim_thompson5910):

we just replace each x with -1 in g(x), then evaluate g(x) = 5x g(-1) = 5(-1) g(-1) = -5

jimthompson5910 (jim_thompson5910):

that will turn f [ h(g(-1)) ] into f [ h(-5) ]

jimthompson5910 (jim_thompson5910):

now find h(-5)

jimthompson5910 (jim_thompson5910):

h(x) = x + 4 h(-5) = -5 + 4 h(-5) = -1

jimthompson5910 (jim_thompson5910):

so f [ h(-5) ] becomes f [ -1 ]

OpenStudy (firejay5):

sorry my allergies are making it hard to see things <--- watery eyes + nose

jimthompson5910 (jim_thompson5910):

last step is to find f(-1) f(x) = x^2 f(-1) = (-1)^2 f(-1) = 1

jimthompson5910 (jim_thompson5910):

so in the end, f [ h(g(-1)) ] = 1 ie [fo(hog)](-1) = 1

OpenStudy (firejay5):

@jim_thompson5910 Thanks so much! :D I actually know how to do it